An assessment of the quality of the electrical power system in Turkey was the topic of an article published in Electrical Engineering​ (Mar. 2013). One measure of quality is the degree to which voltage fluctuations cause light flicker in the system. The perception of light flicker x when the system is set at 380 kV was measured periodically​ (over 10-minute​ intervals). For transformers supplying heavy industry​ plants, the light flicker distribution was found to follow​ (approximately) a normal distribution with μ=2.2​% and σ=0.5​%. Suppose if the perception of light flicker exceeds 2.4​%, the transformer is shut down and the system is reset. How likely is it for a transformer supplying a heavy industry plant to be shut down due to light​ flicker?


1
Expert's answer
2021-07-14T18:05:46-0400

"P(X>2.4)=1-P(z<\frac{2.4-2.2}{0.5})"

"=1-P(z<0.4)"

"=1-0.6554"

"=0.3446"

Let 0.05 be the significance level. 0.3446 is greater than 0.05. Tha chance of electric flicker is 0.3446 which is high. Therefore, it is likely for a transformer supplying a heavy industry plant to be shut down due to light​ flicker.